OpenStudy (anonymous):

The Pie Factory sells an apple pie with a diameter of 16 inches for $10.99. What is the approximate cost per square inch of surface area of the top of the pie? Use 3.14 for p.

jimthompson5910 (jim_thompson5910):

What is the area of the circle?

OpenStudy (anonymous):

it dose not say

jimthompson5910 (jim_thompson5910):

how would you find it

jimthompson5910 (jim_thompson5910):

any ideas?

OpenStudy (anonymous):

multiply ?

jimthompson5910 (jim_thompson5910):

you would use this formula A = pi*r^2

jimthompson5910 (jim_thompson5910):

so because the diameter is 18, the radius is 18/2 = 9

jimthompson5910 (jim_thompson5910):

oh wait 16, not 18

jimthompson5910 (jim_thompson5910):

16/2 = 8 is the radius

OpenStudy (anonymous):

thanks :D

jimthompson5910 (jim_thompson5910):

so A = pi*r^2 A = 3.14*8^2 A = ???

jimthompson5910 (jim_thompson5910):

once you get that area, you divide the cost by the area to get the cost per inch

OpenStudy (anonymous):

thanks :D

jimthompson5910 (jim_thompson5910):

yw, tell me what you get